Interest rate.  Upstate Bank is offering​ long-term certificates of deposit with a face value of ​$100, 000 ​(future value). Bank customers can buy these CDs today for ​$40,000 and will receive the ​$100,000 in 15 years. What interest rate is the bank paying on these​ CDs?
Answer #1

We see that interest rate is given as equal to=(Future Value/Present Value)^(1/t)-1=(100000/40000)^(1/15)-1=6.30%